OpenStudy (anonymous):

Rationalize the denominator of 4/ 5 + square root 2

OpenStudy (anonymous):

\[\frac45 + \sqrt 2 ... or... \frac4{5 + \sqrt2}\]

OpenStudy (anonymous):

the second one

OpenStudy (anonymous):

I mean, which one of my interpretations is correct ? ..

OpenStudy (anonymous):

oh ok!

OpenStudy (anonymous):

:)

OpenStudy (shane_b):

I''m pretty sure she means: \[(\frac{4}{5\sqrt{2}})(\frac{\sqrt{2}}{\sqrt{2}})=\frac{4 \sqrt {2}}{5(2)}=\frac{4 \sqrt {2}}{10}=\frac{2 \sqrt {2}}{5}\]

OpenStudy (anonymous):

\[\frac4{5+\sqrt2} \times \frac{5-\sqrt2}{5-\sqrt2}\] \[= \frac{4(5-\sqrt2)}{5^2 - (\sqrt2)^2}\] Solve the same to get your answer !! :)

OpenStudy (anonymous):

\[20 - 4\sqrt{2} / 23 ?\]

OpenStudy (anonymous):

is that it?

OpenStudy (anonymous):

Absolutely! :)

OpenStudy (anonymous):

thanks for your help! :)

OpenStudy (shane_b):

Did you check the answer?

OpenStudy (anonymous):

yes. it's correct! :)

OpenStudy (shane_b):

Nm....I somehow forgot all about the 5+ sqrt(2)....I just did 5sqrt(2) for the denominator. I was like, I KNOW mine is right!

OpenStudy (anonymous):

Lol its alright :)

OpenStudy (shane_b):

Reading is fundamental :P

OpenStudy (anonymous):

couldn't agree more lolll